# Simple Estimation of Semiparametric Models with Measurement Errors

### v3

@@ Appendix A Regularity Conditions
[…]
@@ Appendix B Proof of Lemma 1
[…]
@@
[…]
@@ Appendix C Proof of Theorem 2
For the true value of the parameter $\alpha_{0}$, we often write $a_{i}\equiv a(\alpha_{0})$, $\overline{a}\equiv\overline{a}(\alpha_{0})$, $a\equiv a(\alpha_{0})$, $a_{i}^{*}\equiv a(\alpha_{0})$, $\overline{a}^{*}\equiv\overline{a}^{*}(\alpha_{0})$, $a^{*}\equiv a^{*}(\alpha_{0})$.
@@ C.1 Auxiliary lemmas
where $\tilde{X}_{i}$ lies in between of $X_{i}^{*}$ and $X_{i}$. Now observe that all the terms following the inequality sign are $o_{p}(1)$. Indeed, this is guaranteed by Markov’s inequality paired with Assumptions 2.1, 2.2, and A.3 (ii). Hence, $\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|=o_{p}(1)$, and we are left to show $\sup_{\theta\in\Theta}\left\|\overline{g}^{*}(\theta)-g^{*}(\theta)\right\|=o_{p}(1)$. This, in turn, follows from the standard ULLN (e.g., Lemma 2.4 in Newey and McFadden, 1994), which also ensures continuity of $g^{*}(\theta)$ on $\Theta$. Hence, we conclude that the assertion of the lemma holds for $g$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ C.2 Proof of Theorem 2
Expanding $\overline{\psi}(\hat{\beta})$ around $\overline{\psi}(\beta_{0n})$ gives
@@
$\overline{\Psi}(\hat{\beta})=\left[\overline{G}(\hat{\theta})-\sum_{k=2}^{K}\hat{\gamma}_{k}\overline{G}_{x}^{(k)}(\hat{\theta}),-\overline{g}_{x}^{(2)}(\hat{\theta}),\dots,-\overline{g}_{x}^{(K)}(\hat{\theta})\right].$
[…]
@@
[…]
@@
Appendix D Details on the sufficient condition (17)
@@
where we used $\mathbb{E}\left[\varepsilon_{Q,i}|X_{i}^{\ast},S_{i},\varepsilon_{i}\right]=0$. Dividing moments $J+2,\ldots,2J+1$ by $\alpha_{1}$ and subtracting them from the moments $12,\ldots,J+$ 1, we obtain

### v4

@@ Appendix A Regularity Conditions
    $\hat{\Xi}\overset{p}{\rightarrow}\Xi$, where $\Xi$ is a symmetric positive definite matrix;
@@ Appendix B Proof of Lemma 1
Since $B_{n}$ is invertible, (B.6) has a unique solution $a_{n}=B_{n}^{-1}c_{n}$. Moreover, $a_{n}$ is bounded since both $B_{n}^{-1}$ and $c_{n}$ are bounded (Assumption 2.1). Hence, we conclude that (B.5) has a unique solution $\gamma_{0n}=\left(\sigma_{n}^{2}a_{2n},\dots,\sigma_{n}^{K}a_{Kn}\right)^{\prime}$. Since (B.5) is satisfied, using Assumption 2.2, we also conclude that
@@
Since $a_{kn}$ is bounded, we conclude that (B.9) holds analogously to (B.8).
@@ Appendix C Proof of Theorem 2
[…]
@@ C.1 Auxiliary lemmas
where $\tilde{X}_{i}$ lies in between of $X_{i}^{*}$ and $X_{i}$. Now observe that all the terms following the inequality sign are $o_{p}(1)$. Indeed, this is guaranteed by Markov’s inequality paired with Assumptions 2.1, 2.2, and A.3(ii). Hence, $\sup_{\theta\in\Theta}\left\|\overline{g}(\theta)-\overline{g}^{*}(\theta)\right\|=o_{p}(1)$, and we are left to show $\sup_{\theta\in\Theta}\left\|\overline{g}^{*}(\theta)-g^{*}(\theta)\right\|=o_{p}(1)$. This, in turn, follows from the standard ULLN (e.g., Lemma 2.4 in NeweyMcFadden1994HB), which also ensures continuity of $g^{*}(\theta)$ on $\Theta$. Hence, we conclude that the assertion of the lemma holds for $g$.
@@
Suppose that the hypotheses of Lemma C.1 are satisfied. Then, $g_{x}^{(k)}\rightarrow g_{x}^{(k)*}$ and $G_{x}^{(k)}\rightarrow G_{x}^{(k)*}$ for $k\in\{0,\dots,K\}$. Suppose also $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$. Then, $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$ and $\overline{G}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}G_{x}^{(k)*}$ for $k\in\{0,\dots,K\}$.
@@
Next, we show $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$. By the triangle inequality,
@@
where the inequality holds with probability approaching one since $\hat{\theta}\in B_{\delta}(\theta_{0})$ with probability approaching one. Note that, By Lemma C.1, $\sup_{\theta\in B_{\delta}(\theta_{0})}\left\|\overline{g}_{x}^{(k)}(\theta)-g_{x}^{(k)*}(\theta)\right\|=o_{p}(1)$ and $\left\|g_{x}^{(k)*}(\hat{\theta})-g_{x}^{(k)*}(\theta_{0})\right\|=o_{p}(1)$, where the second result follows from consistency of $\hat{\theta}$ and continuity of $g_{x}^{(k)*}(\theta)$. Hence, $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$, which completes the proof of the results for $g_{x}^{(k)}$ for all $k\in\{0,\dots,K\}$.
@@
Suppose that the hypotheses of Lemma C.1 are satisfied. Then, under additional Assumptions A.3(iv) and 2.3(ii), we have $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$, $\hat{\gamma}\overset{p}{\rightarrow}0$ and $\hat{\gamma}\overset{p}{\rightarrow}\gamma_{0n}$.
@@
Let $\hat{Q}(\beta)=\overline{\psi}(\beta)^{\prime}\hat{\Xi}\overline{\psi}(\beta)$ and $Q^{*}(\beta)=\psi^{*}(\beta)^{\prime}\Xi\psi^{*}(\beta)$. Notice that (C.1), boundedness of $\psi^{*}(\beta)$, and Assumption A.3(iv) together guarantee that $\sup_{\beta\in\mathcal{B}}\left|\hat{Q}(\beta)-Q^{*}(\beta)\right|=o_{p}(1)$. Next, recall that $\gamma_{0n}\rightarrow 0_{K-1}$ (Lemma 1). Since $\Gamma$ is compact and $\gamma_{0n}\in\Gamma$ (Assumption A.2(i)), $0_{K-1}\in\Gamma$. Consequently, Assumptions 2.3(ii) and A.3(iv) together guarantee that $Q^{*}(\beta)$ is uniquely minimized at $\theta=\theta_{0}$ and $\gamma=0_{K-1}$. Consequently, applying the standard consistency argument (e.g., Theorem 2.1 of NeweyMcFadden1994HB), we conclude that $\hat{\theta}\rightarrow\theta_{0}$ and $\hat{\gamma}\rightarrow 0_{K-1}$. Finally, since $\gamma_{0n}\rightarrow 0$ (Lemma 1), we also have $\hat{\gamma}\overset{p}{\rightarrow}\gamma_{0n}$.
@@
First, note that, by the standard CLT, $n^{-1/2}\sum_{i=1}^{n}g(X_{i}^{*},S_{i},\theta_{0})\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*})$. The rest of the proof is to show that the remaining terms are $o_{p}(1)$. By Assumptions 2.1, 2.2, A.3(ii), Chebyshev’s inequality guarantees
@@ C.2 Proof of Theorem 2
[…]
@@
where $\tilde{\beta}$ lies between $\beta_{0n}$ and $\hat{\beta}$ (and, consequently, $\tilde{\theta}\overset{p}{\rightarrow}\theta_{0}$ and $\tilde{\gamma}\overset{p}{\rightarrow}0$). Next, we argue that $\overline{\Psi}(\hat{\beta})=\Psi^{*}+o_{p}(1)$. Observe
@@
Since $\hat{\theta}\overset{p}{\rightarrow}\theta_{0}$ (Lemma C.3), we can invoke the result of Lemma C.2 to argue that $\overline{g}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}g_{x}^{(k)*}$ and $\overline{G}_{x}^{(k)}(\hat{\theta})\overset{p}{\rightarrow}G_{x}^{(k)*}$ for all $k\in\{0,\dots,K\}$. This, combined with $\hat{\gamma}\rightarrow 0$ (Lemma C.3), ensures that $\overline{\Psi}(\hat{\beta})=\Psi^{*}+o_{p}(1)$ and, analogously, $\overline{\Psi}(\tilde{\beta})=\Psi^{*}+o_{p}(1)$. Coupling these result with Assumption A.3(iv), we conclude that $\overline{\Psi}(\hat{\beta})^{\prime}\hat{\Xi}\overline{\Psi}(\tilde{\beta})\overset{p}{\rightarrow}\Psi^{*\prime}\Xi\Psi^{*}$, which is invertible by Assumption 2.3(i). Hence, (C.3) can be rearranged as (with probability approaching one)
@@
where, by Lemma C.4, $n^{1/2}\overline{\psi}(\beta_{0n})\overset{d}{\rightarrow}N(0,\Omega_{gg}^{*})$. Hence, we conclude
$n^{1/2}(\hat{\beta}-\beta_{0n})\overset{d}{\rightarrow}N(0,\Sigma^{*}),$
@@
Appendix D Proof of Lemma 3
As in the previous proofs, we use notation $\sigma_{n}^{2}\equiv\mathbb{E}[\varepsilon_{i}^{2}]\equiv\sigma_{\varepsilon}^{2}$, $\hat{\sigma}^{2}_{n}\equiv\hat{\sigma}_{\varepsilon}^{2}$, and $\hat{\sigma}_{n}^{K}\equiv\hat{\sigma}_{\varepsilon}^{K}$.
First, we show that (19) holds. By Theorem 2, we have $\hat{\sigma}_{n}^{2}=\sigma_{n}^{2}+O_{p}(n^{-1/2})$. Since $K$ is even, (19) can be obtained by expanding
$\hat{\sigma}_{n}^{K}$ $=(\sigma_{n}^{2}+O_{p}(n^{-1/2}))^{K/2}=\sum_{\ell=0}^{K/2}\binom{n}{\ell}(\sigma_{n}^{2})^{K/2-\ell}(O_{p}(n^{-1/2}))^{\ell}$
$=\sigma_{n}^{K}+\sum_{\ell=1}^{K/2}\binom{n}{\ell}(\sigma_{n}^{2})^{K/2-\ell}O_{p}(n^{-\ell/2})$
$=\sigma_{n}^{K}+O_{p}(\max\{\sigma_{n}^{K-2}n^{-1/2},n^{-K/4}\}).$
Next, we prove the first statement of the lemma. First, using Lemma C.2, we have $\overline{g}_{x}^{(K)}(\hat{\theta})=g_{x}^{(K)*}+o_{p}(1)$. Here, $\|g_{x}^{(K)*}\|$ is bounded away from zero and above, with the former implied by Assumption 2.3(i). Likewise, since the Jacobian corresponding to $\hat{\beta}_{L}$ (involved in the construction of $B$) is a submatrix of the Jacobian corresponding to $\hat{\beta}_{K}$, we conclude that $\|Bg_{x}^{(K)*}\|$ is also bounded from zero and above. Hence, using consistency of $\hat{B}$ and $\hat{\Sigma}$, we conclude that, with probability approaching one, for some $\underline{C},\overline{C}>0$, we have
$\underline{C}\sqrt{n}\hat{\sigma}_{n}^{K}\leq\max_{1\leq\ell\leq\dim(\beta)}|\hat{\Sigma}_{\ell\ell}^{-1/2}\sqrt{n}\hat{\sigma}^{K}\hat{B}_{\ell\cdot}\overline{g}_{x}^{(K)}(\hat{\theta})|\leq\overline{C}\sqrt{n}\hat{\sigma}_{n}^{K}.$ (D.1)
To prove the first part of the lemma it is sufficient to show that $\delta_{n}\sigma_{n}^{L+1}=o_{p}(n^{-1/2})$. Let $n_{j}$ index a subsequence of $\delta_{n_{j}}$ such that $\delta_{n_{j}}=1$. If this subsequence is finite, the statement is trivial. Otherwise, we need to show that $\sigma_{n_{j}}^{L+1}=o(n_{j}^{-1/2})$ along this subsequence. We prove this by contradiction. Suppose $\sigma_{n_{j}}^{L+1}=o(n_{j}^{-1/2})$ does not hold meaning that there exists another subsequence within $n_{j}$, denoted by $n_{m}$ such that for all sufficiently large $m$ we have $\sqrt{n_{m}}\sigma_{n_{m}}^{L+1}\geq C>0$ and $\delta_{n_{m}}=1$. Since we have $\sigma_{n_{m}}\geq Cn_{m}^{-\frac{1}{2(L+1)}}$ and $\sqrt{n_{m}}\sigma_{n_{m}}^{K}\geq Cn_{m}^{-(K-L-1)/2(L+1)}$, using (19), we conclude that, for some $C>0$, with probability approaching one, we have
$\sqrt{n_{m}}\hat{\sigma}_{n_{m}}^{K}=\sqrt{n_{m}}\sigma_{n_{m}}^{K}+O_{p}\left(\sigma_{n_{m}}^{K-2}+n_{n_{m}}^{-\left(K-2\right)/4}\right)\geq Cn_{m}^{-(K-L-1)/2(L+1)}$ (D.2)
because the remainder $O_{p}(\sigma_{n_{m}}^{K-2}+n_{n_{m}}^{-\left(K-2\right)/4})$ is dominated by $\sqrt{n_{m}}\sigma_{n_{m}}^{K}$ in this case. At the same time, $\delta_{n_{m}}=1$ together with (D.1) implies that $\sqrt{n_{m}}\hat{\sigma}_{n_{m}}^{K}\leq C\varkappa_{n_{m}}$ for some $C>0$. This contradicts (D.2) since $\varkappa_{n}n^{\left(K-L-1\right)/\left(2L+2\right)}\rightarrow 0$. The contradiction completes the proof of the first part.
We now turn to the second part of the lemma. First, note that, using Jensen’s inequality, we obtain
$\sqrt{n}\hat{\sigma}_{n}^{K}=\sqrt{n}(\sigma_{n}^{2}+O_{p}(n^{-1/2}))^{K/2}\leq C\sqrt{n}\sigma_{n}^{K}+O_{p}(n^{-(K-2)/4}).$
In particular, for $\varkappa_{n}=n^{-\left(K-L-1\right)/\left(2L+2\right)}\left(\ln n\right)^{-a}$ and $\sigma_{n}=o(n^{-\frac{1}{2L+2}-\epsilon})$, the above implies that $\sqrt{n}\hat{\sigma}_{n}^{K}/\varkappa_{n}=o_{p}(1)$. This, together with (D.1), implies that $\delta_{n}=1$ with probability approaching one, which completes the proof.
Appendix E Details on the sufficient condition (17)
@@
where we used $\mathbb{E}\left[\varepsilon_{Q,i}|X_{i}^{\ast},S_{i},\varepsilon_{i}\right]=0$. Dividing moments $J+2,\ldots,2J+1$ by $\alpha_{1}$ and subtracting them from the moments $2,\ldots,J+1$, we obtain